Prince​ Electronics, a manufacturer of consumer electronic​ goods, has five distribution centers in different regions of the country. For one of its​ products, a highspeed modem priced at ​$330 per​ unit, the average weekly demand at each distribution center is 85 units. Average shipment size to each distribution center is 450 ​units, and average lead time for delivery is 3 weeks. Each distribution center carries 3 ​weeks' supply as safety stock but holds no anticipation inventory. a. On​ average, how many dollars of pipeline inventory will be in transit to each distribution​ center? ​$84150 . ​(Enter your response as an​ integer.) b. How much total inventory​ (cycle, safety, and​ pipeline) does Prince hold for all five distribution​ centers? nothing units. ​(Enter your response as an​ integer.)
Answer #1

a.
Value of pipeline inventory = Average weekly demand*weeks supply as safety stock*price per unit = 85*3*330 = 84150
b.

Cycle inventory = no of distribution center*(Average shipment size/2) = (450/2) = 225

Safety inventory = no of weeks*average weekly demand = 3*85 = 255

Pipeline inventory = lead time*average weekly demand = 3*85 = 255

Total inventory = Cycle inventory+Safety inventory +Pipeline inventory = 5*(225+255+255) = 3675 units